Could someone Please help me!!!
It would be most appreciated:)
Question:
1)(3+sqrt(-16))(4+sqrt(-9)) This question I'm stuck half way:(
&
2)(-3+4i)^2 On this question I got 9+16i but I don't think its right???
Thank you!!!



Answer by stanbon(75887)

1)(3+sqrt(-16))(4+sqrt(-9))
----------------
= (3+4i)(4+3i)
-----
= 12+ 9i + 16i +12i^2
----
= 12 +25i -12
---
= 25i
==============
2)(-3+4i)^2
---
= (-3+4i)(-3+4i)
---
= 9 -12i -12i +16i^2
-----
= 9 -24i -16
---
= -7-24i
================
Cheers,
Stan H.